OpenStudy (anonymous):

4+(sqaure root 3/2)^2

OpenStudy (anonymous):

lol.. uh.. here we go again.... :)

OpenStudy (anonymous):

let me clarify the question first lol

OpenStudy (anonymous):

\[4 + (\frac{\sqrt{3}}{2})^2 \; \; ??\]

OpenStudy (anonymous):

its sec^2 pi/3 +sin^2 pi/3 :)

OpenStudy (anonymous):

\[(\frac{\sqrt{3}}{2})^2 = ??\]

OpenStudy (anonymous):

..what ?

OpenStudy (anonymous):

Do you know the value of this??

OpenStudy (anonymous):

would it be 3/4 :) ?

OpenStudy (anonymous):

\[(\frac{a}{b})^2 = \frac{a^2}{b^2}\]

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

\[4+ \frac{3}{4} = ??\] Can you solve this??

OpenStudy (anonymous):

Firstly you have to make the denominator same... Go ahead, if any doubt, then let us know..

OpenStudy (anonymous):

i got 4 3/4

Parth (parthkohli):

Yes, correct.

OpenStudy (anonymous):

but my text boom says 4 1/4

OpenStudy (anonymous):

thats why i got stuck

OpenStudy (anonymous):

In Mixed fraction form, you are right..

OpenStudy (anonymous):

Oh My God.. You and your book...

OpenStudy (anonymous):

Are you sure that is sin^2(pi/3) and not sin^2(pi/6) ??

OpenStudy (anonymous):

Sorry pi/6.. Ha ha ha..

OpenStudy (anonymous):

haha ur right !!

OpenStudy (anonymous):

@yasmin95 , there is no hurry.. Please post the right question... Anyways, well done..